2.A stone moves along the path x=t^3+1,y=t^3, z=2t+5, where t denote time. What is the component of (dp)/dt ?
3i+2j-2k
3i-2j+2k
3i+2j+2k
-3i+2j+2k
1
Expert's answer
2020-05-19T20:16:18-0400

A stone moves along the path "x=t^3+1, y=t^2, z=2t+5"


"dx\/dt=3t^2, dy\/dt=2t, dz\/dt=2"

If "t=1>0, dp\/dt=(3, 2, 2)"

3i+2j+2k